in a certain statistics class, the marks obtained by students on a class test followed a normal distribution with a mean of 68 and a standard deviation of 10. What is the probability that the mean test mark from a sample of 25 students from the class was more than 70?
Answer by stanbon(57962)

in a certain statistics class, the marks obtained by students on a class test followed a normal distribution with a mean of 68 and a standard deviation of 10. What is the probability that the mean test mark from a sample of 25 students from the class was more than 70?
---
t(70) = (70-68)/[10/sqrt(25)] = 1
---
P(x-bar > 70) = P(t > 1 when df = 24) = 0.1636
=====
Cheers,
Stan H.